# A Novel Approach to Reduce Derivative Costs in Variational Quantum Algorithms

### v2

@@ III.1 Cost analysis and comparison between DM and QNDM approaches
To make a comparison between the two approaches, we use the cost $\mathcal{C}\left(g_{j}^{i}\right)$ ($i=QNDM,\,DM$) that gives the scaling in terms of the number of gates required to compute a derivative in the cost function for the two algorithms represented in Figs. 2 and 3. In addition, we use the resource ratio $\mathcal{C}\left(g_{j}^{DM}/g_{j}^{QNDM}\right)$.
The resource costs are functions of the number of Pauli strings $J$, the number of qubits $n$, and the number of logical operators $k$ used in the circuits to calculate a single derivative (where in $k$ we include both single-qubit and two-qubit gates). As in the previous section, $N_{DM}$ and $N_{QNDM}$ represent the number of shots or repetitions for DM and QNDM, respectively. The values extracted from Refs. Solinas et al. (2023); Cerezo et al. (2021); Mari et al. (2021) are detailed in Tab. 1.
It is important to note that we assign two different values for the required number of shots because, as discussed in Sec. III, the MSE in the case of DM depends linearly on the number of Pauli strings $J$. Therefore, to obtain the same MSE error in the derivative estimation for QNDM and DM, the number of shots for DM is $N_{DM}\propto N_{QNDM}J$ [see Eq. (28)].
@@
DM $N_{DM}$ $\mathcal{C}\left(g_{j}^{DM}\right)=2N_{DM}J(k+n)$
QNDM $N_{QNDM}$ $\mathcal{C}\left(g_{j}^{QNDM}\right)=N_{QNDM}(3k+8Jn)$
Table 1: Resource cost function to compute first-order derivatives with the DM and the QNDM approaches.
Given these premises, we are interested in two regimes. First, when $k\gg nJ$, i.e., the number of logical gates to implement $U(\vec{\theta})$ exceeds the number of Pauli strings $J$ (time the number of qubits). From Tab. 1 and Ref. Solinas et al. (2023), we have that the resource ratio scales as
$\mathcal{C}\left(\frac{g_{j}^{DM}}{g_{j}^{QNDM}}\right)=\frac{2J}{3}\frac{2\sigma_{s}^{2}\lambda^{2}(1-(2P_{0}-1)^{2})}{\sigma^{2}_{D}}\sum_{i=1}^{J}h^{2}_{i}=\mathcal{O}(J^{2}).$ (29)
This behavior shows that the cost to calculate a derivative with DM is $J^{2}$ higher than QNDM so that, in this regime, we have quadratic speed up with the number of Pauli strings. This arises from two components: firstly, the inherent linear advantage that QNDM  holds over DM  Solinas et al. (2023), and secondly, the dependency of the MSEDM on the number of $J$, which is absent in MSEQNDM. This regime is particularly relevant for quantum chemistry simulations. For instance, in simulations of moderately complex molecules Wecker et al. (2015), we have $k\approx 10^{9}-10^{10}$ operations to implement $U(\vec{\theta})$, $n\approx 10^{2}-10^{3}$ (the expected dimension of quantum computers in the next years), and $J>10^{3}$ as the number of Pauli strings in the Hamiltonian. With these numbers, $k\gg nJ$, resulting in a reduction of roughly $J^{2}$ resources when using the QNDM approach. This advantage is likely to increase for larger and more complex molecules.
@@
$\mathcal{C}\left(\frac{g_{j}^{DM}}{g_{j}^{QNDM}}\right)=\frac{k}{4n}\frac{2\sigma_{s}^{2}\lambda^{2}(1-(2P_{0}-1)^{2})}{\sigma^{2}_{D}}\sum_{i}h^{2}_{i}=\mathcal{O}(Jk).$ (30)
Whereas the $k$ factor comes from the linear cost advantage of QNDM with respect to DM for this regime (see Tab. 2), the $J$ remains as a consequence of the different dependency of the MSEDM and MSEQNDM. This scenario is promising for various applications in VQAs as pointed out in Flynn et al. (2022).
@@ IV.1 Errors Analysis
In this section, we present a numerical comparison between errors estimated using the two methods. For the numerical analysis, we have used the simulator provided by IBM Quantum known as Aer IBM Quantum (2024). As noticed at the end of Sec. III, we expect an advantage from the QNDM approach (with respect to the DM approach).
The value of the derivatives $g_{j}$ depends on many parameters:
  1. (i)
@@
  3. (iii)
    the direction $j$ along which is calculated the derivative;
  4. (iv)
    the Pauli strings $\hat{P}_{j}$ that appear in $\hat{M}$;
@@
For this reason, to ensure statistical significance, we need to average over these parameters. In the following, we use the term “realization” to denote a random choice of all the relevant parameters and possibilities described in the points from (i) to (v).
Figure 4: The main plot represents the value of the derivative over the $j$-th direction $\mu(g_{j}^{i})$ [$i=$DM (red markers) and $i=$QNDM (blue markers)], averaged over a collection of different realizations, and shown as a function of the number of Pauli strings. In the inset, we show the ratio between the $MSE_{DM}/MSE_{QNDM}$. The simulated quantum system has $n=10$ qubits and the number of shots is fixed at $N=10^{4}$. Each point is averaged for $L=100$ realizations implemented as discussed in the main text.
In Fig. 4 we plot the averaged derivative $\mu(g_{j}^{i})$ over a randomly chosen direction $j$, and their errors as a function of the number of Pauli strings $J$. The DM  and QNDM  results correspond to the red and blue markers, respectively. In this case, the average of $\mu(g_{j}^{i})$ is performed over a collection of different realizations averaged over the above parameter (i) to (iv). For each point, we randomly selected $L=100$ realizations. Also, the errors associated with the points are the result of the average of the errors of the corresponding realizations.
Notice that, while the values obtained with the two protocols are compatible within the error bars, the uncertainty in the DM approach increases with $J$ while the QNDM does not. This means that, with the DM approach we must increase the number of shots to reach the same precision obtained with the QNDM one. This is further exemplified by the inset in Fig. 4 where we plot the ratio of the MSEs. As can be seen, the MSEDM is a factor of $10$ larger than MSEQNDM already for $J=100$, and the difference grows linearly with $J$.
@@ IV.2 Costs Comparison
For practical situations, it is of paramount importance to quantitatively estimate any advantage of the QNDM algorithm over DMand in which conditions this is reached. From the discussion in Sec. III.1, we expect the QNDM approach to perform better, i.e., it needs fewer resources, in some specific but interesting regimes. Aiming in this direction, we present a full numerical analysis of the resources needed to calculate the derivative of the cost function.
To have a meaningful comparison, we have fixed the precision, i.e., the MSE required for the derivative evaluation for both approaches. Consequently, for the DM approach, we must increase the number of shots, as we already showed in Eq. (28) and Fig. 4.
@@
Figure 5: Regime $k\gg nJ$. $a)$ The averaged number of resources $\mathcal{C}(g_{j}^{i})$ needed to estimate the derivative of the cost function along the direction $j$; here, $i=QNDM$ (orange line) and $i=DM$ (blue line). $\mathcal{C}(g_{j}^{i})$ is plotted as a function of the number of logical operators $k$ for a fixed $nJ=240$. $b)$ Ratio between DM and QNDM resources numbers $\left(\mathcal{C}\left(g_{j}^{DM}/g_{j}^{QNDM}\right)\right)$ as a function of Pauli string $J$ each colored line corresponds at a fixed $k$. The quantum circuit is composed of $n=10$ qubits and the average is taken over $L=50$ realization as discussed in the main text. The number of shots for the QNDMis $N_{QNDM}=500$, for DM, it is calculated for each realization by (28), such that the MSE errors are equal for both methods.
@@
The results of Fig. 5 illustrate the expected advantage of QNDM over the DM approach. Panel a), on the left, shows the resource cost $\mathcal{C}\left(g_{j}^{i}\right)$ for the two approaches in the limit $k\gg nJ$. They are plotted as a function of the number of logical operators $k$ needed to implement $U(\vec{\theta})$. The simulations have been done for $n=10$ qubits and a with $J=24$ different Pauli strings. Each value in Fig. 5 is calculated by averaging over different realizations with randomly chosen values of the parameters $J$, $k$, and the set of Pauli strings in $\hat{M}$. The average is performed also over the unitary transformations $U(\vec{\theta})$. This is obtained through a series of layers parameterized by $U_{j}(\vec{\theta}_{j})$ and entanglement layers denoted as $V_{j}$, as illustrated in Eqs. (3) and (4). Within the parameterized layer, the constituent single-qubit rotations $R^{j}_{i}(\theta_{i}^{j})$, and their corresponding parameters $\vec{\theta}_{i}^{j}$, are selected at random.
@@
In panel a) of Fig. 5, as expected, QNDM needs fewer resources than DM  to evaluate the derivatives. Less obvious is that already for a limited number of logical operators, i.e. $k\sim 10^{3}$, we save more than $10^{2}$ logical gates.
@@
In panel b) of Fig. 5, on the right, we show the ratio of the resources in terms of the number of Pauli strings $J$, comparing different possible choices for $k$. Already for $k\sim 10^{3}$, the ratio approaches the theoretical prediction in Eq. (29) (shown as a dashed black curve in the plot) and the DM method needs more than $100$ times more resources than the QNDM one. Due to the increasing number of shots needed for the DM when $J$ increases, we expect it to reach already $10^{6}$ for simple practical applications Wecker et al. (2015).
@@
Figure 6: Regime $k\ll nJ$. The averaged number of resources $\mathcal{C}(g_{j}^{i})$ needed to estimate the derivative of the cost function along the direction $j$ ($i=QNDM$(orange line) and $i=DM$(blue line)). $\mathcal{C}(g_{j}^{i})$ is plotted as a function of the product $nJ$ for a fixed number of logical operators $k=0.5\times 10^{3}$. $b)$ Ratio between DM and QNDM resources $\left(\mathcal{C}\left(g_{j}^{DM}/g_{j}^{QNDM}\right)\right)$ as a function of the number of logical operators $k$ each colored line corresponds at a fixed $nJ$. The quantum circuit is composed of $n=10$ qubits and the average is taken over $L=50$ realization as discussed in the main text. The number of shots for the QNDMis $N_{QNDM}=500$, for DM, it is calculated for each realization by (28), such that the MSE errors are equal for both methods.
@@
The regime in which $k\ll nJ$ is analyzed in Fig. 6. Panel a), on the left, shows the resource cost function $\mathcal{C}(g_{j}^{i})$ ($i=QNDM,DM$) obtained to estimate the derivative along the direction $j$. The values are plotted as a function $nJ$. The simulations are done for $n=10$ qubits with $k=0.5\cdot 10^{3}$ fixed. Each value in Fig. 6 is calculated by performing an average over a collection of different realizations, as discussed above. As delineated in tab. 1, for DM (the blue line) the number of resources increases quadratically in the number of $J$. Already for $kJ\sim 10^{3}$, the DM method need $10^{4}$ more resources than QNDM one.
@@
In panel b), on the right, we show the ratio of the resources as a function of the number of logical operators $k$ and for different values of $nJ$. The expected advantage of the QNDM  approach is confirmed in this regime. Already for $nJ\sim 10^{3}$ the DM approach needs almost $10^{5}$ times more resources than the QNDM one.
@@
Notice that the results obtained complement the analytic analysis provided in Solinas et al. (2023). Specifically, the error analysis, Sec. III and Sec. IV.1, shows that, for the QNDM, the statistical uncertainty, i.e., the mean square error, is independent of the number of Pauli strings $J$ appearing in the quantum observable, whereas for the DM, it increases linearly with $J$. This ensures additional savings in resources, as shown in Fig. 5 and Fig. 6, compared to what was anticipated in our analytical analysis Solinas et al. (2023). Additionally, in all our simulations and both regimes, we have considered relatively small and simple systems to provide a more conservative estimate. We might expect a further increasing advantage in realistic computational problems, as they usually require more logical gates and involve more complex observables $\hat{M}$.

### v3

@@ III.1 Cost analysis and comparison between DM and QNDM approaches
To make a comparison between the two approaches, we use the cost $\mathcal{C}\left(g_{l}^{i}\right)$ ($i=QNDM,\,DM$) that gives the scaling in terms of the total number of gates required to compute a derivative in the cost function for the two algorithms represented in Figs. 2 and 3. In addition, we use the resource ratio $\mathcal{C}\left(g_{l}^{DM}/g_{l}^{QNDM}\right)=\mathcal{C}\left(g_{l}^{DM}\right)/\mathcal{C}\left(g_{l}^{QNDM}\right)$.
The resource costs express the number of gates, both single-qubit and two-qubit gates, used in one of two approaches. They are functions of the number of Pauli strings $J$, the number of qubits $n$, and the number of logical operators $k$ used in the unitary transformation $U(\theta)$, Eq. (1). As in the previous section, $N_{DM}$ and $N_{QNDM}$ represent the number of shots or repetitions for the two approaches.
For DM approach, described in sec. II.2, a circuit with $n$ qubits is transformed by a unitary operator $U(\theta-se_{l})$, and $P_{i}$ is measured $N_{DM}$ times. This process is repeated for each of the $J$ Pauli strings in $\hat{M}$ to calculate the expectation value $\langle M(\theta-se_{l})\rangle$. The same procedure is then repeated with the unitary operator $U(\theta+se_{l})$ to obtain $\langle M(\theta+se_{l})\rangle$. Summing all the contributions we get a computational cost equal to $2N_{DM}J(k+n)$. For QNDM approach, described in sec. II.3, a circuit with $n+1$ qubits is initially transformed by the operator $U_{1}=U(\theta-se_{l})$, followed by the application of the exponential operator $U_{-}=\exp\{-i\lambda Z_{a}\otimes\hat{M}\}$. Next, the operator $U_{2}=U^{\dagger}(\theta-se_{l})U(\theta-se_{l})$ is applied. Finally, the second exponential operator $U_{+}=\exp\{+i\lambda Z_{a}\otimes\hat{M}\}$ is applied, and the detector phase is measured. This procedure is repeated $N_{QNDM}$ times. The resulting computational cost is $N_{QNDM}(3k+8Jn)$. The computed values of the resource costs are summarized in Tab. 1.
@@
DM $N_{DM}$ $\mathcal{C}\left(g_{l}^{DM}\right)=2N_{DM}J(k+n)$
QNDM $N_{QNDM}$ $\mathcal{C}\left(g_{l}^{QNDM}\right)=N_{QNDM}(3k+8Jn)$
Table 1: Resource cost function to compute first-order derivatives with the DM and the QNDM approaches. As can be seen, the QNDM has a clear cost advantage. However is import to emphasize that in the QNDM approach we run a more deeper quantum circuit, Fig. 3 , than DM, Fig. 2.
In this analysis, we focus on two regimes. The first occurs when $k\gg nJ$, meaning that the number of logical gates required to implement $U(\vec{\theta})$ exceeds the number of Pauli strings $J$ (multiplied by the number of qubits). From Tab. 1 and Ref. Solinas et al. (2023), we have that the resource ratio scales as
$\mathcal{C}\left(\frac{g_{l}^{DM}}{g_{l}^{QNDM}}\right)=\frac{2J}{3k}\frac{N_{DM}}{N_{QNDM}}=\mathcal{O}(J).$ (29)
This result indicates that the cost of computing a derivative using DM is $J$ times higher than QNDM , leading to a linear speedup in this regime as the number of Pauli strings increases. This regime is particularly relevant for quantum chemistry simulations. For instance, in simulations of moderately complex molecules Wecker et al. (2015), typical values are $k\approx 10^{9}-10^{10}$ operations for implementing $U(\vec{\theta})$, $n\approx 10^{2}-10^{3}$ (corresponding to the expected size of near-term quantum processors), and $J>10^{3}$ as the number of Pauli strings in the Hamiltonian. Given these parameters, the condition $k\gg nJ$ holds, resulting in a reduction of approximately $J$ in resource usage when employing the QNDM approach. This advantage is expected to grow as simulations scale to larger and more complex molecules.
@@
$\mathcal{C}\left(\frac{g_{l}^{DM}}{g_{l}^{QNDM}}\right)=\frac{k}{4n}\frac{N_{DM}}{N_{QNDM}}=\mathcal{O}(k).$ (30)
Whereas the $k$ factor comes from the linear cost advantage of QNDM with respect to DM for this regime (see Tab. 2). This scenario is promising for various applications in VQAs as pointed out in Flynn et al. (2022).
@@ IV.1 Errors Analysis
In this section, we present a numerical comparison of error estimates obtained using the two methods. To gain a comprehensive understanding of the errors, we conducted multiple tests across different configurations. Here, we present one such test, where the number of qubits is $n=10$ and the Pauli strings in $\hat{M}$ are randomly selected from the set of all possible Pauli string operators. The coefficients $h_{i}$ in Eq. (6) are drawn from a Gaussian distribution centered at 00 with varying standard deviations. In all simulations, the results remain consistent with those obtained for a standard deviation of $5$, which we report here. The analysis is performed as a function of the number of Pauli strings $J$. As discussed in Sec. III, to approximate the derivative of the $\mathcal{G}_{\lambda}$, the value of $\lambda$ must be taken small enough to justify the linear approximation in Eq. (13). As a consequence, the value of $\lambda$ depends (in a complex way) on the coefficients of the observable $\hat{M}$. For this case, we set a value of $\lambda$ equal to the square root of the inverse of the sum of the absolute values of all coefficients $h_{i}$, $\lambda=1/\sqrt{\sum_{i}|h_{i}|}$. For the numerical analysis, we have used the simulator provided by IBM Quantum known as Aer IBM Quantum (2024). All the simulations are run to a shot count of $500$, a standard value for optimization.
The value of the derivatives $g_{l}$ depends on many parameters:
  1. (j)
@@
  3. (jjj)
    the direction $l$ along which is calculated the derivative;
  4. (jv)
    the Pauli strings $\hat{P}_{i}$ that appear in $\hat{M}$;
@@
For this reason, to ensure statistical significance, we need to average over these parameters. In the following, we use the term “realization” to denote a random choice of all the relevant parameters and possibilities described in the points from $(j)$ to $(jv)$.
Figure 4: The plot represents the average derivative $\mu(g_{l}^{i}(J))$ along with the average mean squared error $\mu(MSE^{i}(J))$ [$i=$DM (red markers) and $i=$QNDM (blue markers)]. Each point is a function of the number of Pauli strings $J$. The results correspond to a shot count of $500$. The simulated quantum system consists of $n=10$ qubits. Each point is averaged over $L=100$ realizations, as discussed in the main text.
In Fig. 4, we plot the derivative average $\mu(g_{l}^{i}(J))$ on the vertical axis, along with the average mean squared error $\mu(MSE^{i}(J))$. For each point, we calculate $L=100$ derivatives $g_{l}^{i}(J)$, varying the above parameters $(j)$ to $(v)$, and calculate their corresponding $MSE(g_{l}^{i}(J))$. Finally, we average all the derivatives and their corresponding MSEs.
$\mu(g_{l}^{i}(J))=\frac{1}{L}\sum^{L=100}_{\Omega}g_{l}^{i}(J),\quad\mu(MSE^{i}(J))=\frac{1}{L}\sum^{L=100}_{\Omega}MSE(g_{l}^{i}(J)),$ (31)
where the sum over $\Omega$ is intended over all the indices $(j),(jj),(jjj),(jv),(v)$. The horizontal axis represents the number of Pauli strings $J$. The DM and QNDM results correspond to the red and blue markers, respectively.
Notice that although the ratio between the errors of the DM and QNDM approaches remains constant as the number of Pauli strings increases, the QNDM approach with the same number of shots provides a growing advantage in terms of $MSE$. This means that, with the DM approach we must increase the number of shots to reach the same precision obtained with the QNDM one.
@@ IV.2 Costs Comparison
For practical situations, it is of paramount importance to quantitatively estimate any advantage of the QNDM algorithm over DM and in which conditions this is reached. From the discussion in Sec. III.1, we expect the QNDM approach to perform better, i.e., it needs fewer resources, in some specific but interesting regimes. See for example Eqs. (29) and (30). Aiming in this direction, we present a full numerical analysis of the resources needed to calculate the derivatives of the cost function. For a meaningful comparison, we fix the number of shots for the QNDM approach is set to $N_{QNDM}=500$, while for the DM approach, the number of shots is adjusted for each instance to ensure $MSE_{QNDM}=MSE_{DM}$. This means that if the MSE of QNDM for 500 shots is lower than the MSE of DM with 500 shots, the number of shots for DM must be increased (or vice versa) to satisfy the condition $MSE_{QNDM}=MSE_{DM}$. As in the previous section, we present our cost analysis for a general example in which all coefficients of the observable $\hat{M}$ are drawn from a Gaussian distribution centered at 00 with a standard deviation of $5$. As before, we took a value of $\lambda$ equal to the square root of the inverse of the sum of the absolute values of all coefficients $h_{i}$, $\lambda=1/\sqrt{\sum_{i}|h_{i}|}$.
@@
Figure 5: Regime $k\gg nJ$. $a)$ The averaged number of resources $\mathcal{C}(g_{l}^{i})$ needed to estimate the derivative of the cost function along the direction $l$; here, $i=QNDM$ (orange line) and $i=DM$ (blue line). $\mathcal{C}(g_{l}^{i})$ is plotted as a function of the number of logical operators $k$ for a fixed $nJ=240$. $b)$ Ratio between DM and QNDM resources numbers $\left(\mathcal{C}\left(g_{l}^{DM}/g_{l}^{QNDM}\right)\right)$ as a function of Pauli string $J$ each colored line corresponds at a fixed $k$. The quantum circuit is composed of $n=10$ qubits and the average is taken over $L=50$ realization as discussed in the main text. The number of shots for the QNDM is $N_{QNDM}=500$, for DM, it is calculated for each realization such that the MSE errors are equal for both methods.
@@
The results of Fig. 5 illustrate the expected advantage of QNDM over the DM approach. Panel a), on the left, shows the resource cost $\mathcal{C}\left(g_{l}^{i}\right)$ for the two approaches in the limit $k\gg nJ$. They are plotted as a function of the number of logical operators $k$ needed to implement $U(\vec{\theta})$. The simulations have been done for $n=10$ qubits and a with $J=24$ different Pauli strings. Each value in Fig. 5 is calculated by averaging over different realizations with randomly chosen values of the parameters $J$, $k$, and the set of Pauli strings in $\hat{M}$. The average is performed also over the unitary transformations $U(\vec{\theta})$. This is obtained through a series of layers parameterized by $U_{j}(\vec{\theta}_{j})$ and entanglement layers denoted as $V_{j}$, as illustrated in Eqs. (3) and (4). Within the parameterized layer, the constituent single-qubit rotations $R^{j}_{i}(\theta_{i}^{j})$, and their corresponding parameters $\vec{\theta}_{i}^{j}$, are selected at random.
@@
In panel a) of Fig. 5, as expected, QNDM needs fewer resources than DM to evaluate the derivatives. Less obvious is that already for a limited number of logical operators, i.e. $k\sim 10^{2}$, we achieve a reduction of more than an order of magnitude in the number of logical gates.
@@
In panel b) of Fig. 5, on the right, we show the ratio of the resources in terms of the number of Pauli strings $J$, comparing different possible choices for $k$. Already for $k\sim 10^{2}$, the ratio closely follows the theoretical linearity predicted in Eq. (29). In this regime, the DM method requires more than 35 times the resources of the QNDM approach.
@@
Figure 6: Regime $k\ll nJ$. The averaged number of resources $\mathcal{C}(g_{j}^{i})$ needed to estimate the derivative of the cost function along the direction $j$ ($i=QNDM$(orange line) and $i=DM$(blue line)). $\mathcal{C}(g_{j}^{i})$ is plotted as a function of the product $nJ$ for a fixed number of logical operators $k=0.5\times 10^{3}$. $b)$ Ratio between DM and QNDM resources $\left(\mathcal{C}\left(g_{j}^{DM}/g_{j}^{QNDM}\right)\right)$ as a function of the number of logical operators $k$ each colored line corresponds at a fixed $nJ$. The quantum circuit is composed of $n=10$ qubits and the average is taken over $L=50$ realization as discussed in the main text. The number of shots for the QNDMis $N_{QNDM}=500$, for DM, it is calculated for each realization such that the MSE errors are equal for both methods.
@@
The regime in which $k\ll nJ$ is analyzed in Fig. 6. Panel a), on the left, shows the resource cost function $\mathcal{C}(g_{l}^{i})$ ($i=QNDM,DM$) obtained to estimate the derivative along the direction $l$. The values are plotted as a function $nJ$. The simulations are done for $n=10$ qubits with $k=0.5\cdot 10^{3}$ fixed. Each value in Fig. 6 is calculated by performing an average over a collection of different realizations, as discussed above. As delineated in tab. 1, for DM (the blue line) the number of resources increases linearly in the number of $J$. Already for $nJ\sim 10^{3}$, the DM method already requires 10 times more resources than the QNDM one.
@@
In panel b), on the right, we show the ratio of the resources as a function of the number of logical operators $k$ and for different values of $nJ$. The expected advantage, Eq. (30), of the QNDM approach is confirmed in this regime. Already for $nJ\sim 10^{3}$ the DM approach needs $40$ times more resources than the QNDM one.
@@
Notice in all our simulations and both regimes, we have considered relatively small and simple systems to provide a more conservative estimate. We might expect a further increasing advantage in realistic computational problems, as they usually require more logical gates and involve more complex observables $\hat{M}$.